# with theoretical and observational applications

### v1

@@ 4.3.1 Analytic solution $a(t)$
We turn to the radiation - matter phase that lasts between $N_{r}$ and $N_{m}$, as defined in (3.17). The phase starts with radiation domination and finishes with matter domination, in agreement with the scaling in $a$ of $\rho_{r}$ and $\rho_{m}$. We thus expect $P_{r}$ and $P_{m}$ to provide asymptotic expressions of a complete analytical solution for this whole phase, at small and large times respectively. We are interested in finding such a complete solution, interpolating between these two solutions.
@@
where $t_{*}$ captures the possible integration constant. This solution has been known for a long time: see [30, 29], and [31] with curvature, or [32, App. B] for a recent account.
@@
Figure 15: $X(T)$, standing up to normalisation for $a(t)$, as the analytical solution for the radiation - matter phase. The three colors of the curve correspond to the three analytical parts of the solution given in (4.24). The continuity of the solution is manifest.
@@
The analytical solution (4.24) should in very good approximation reproduce $a(t)$ during the radiation - matter phase, and in particular, reproduce the $\Lambda$CDM solution from $t=0$ until matter domination; in the later phase of dark energy domination, the analytical solution should differ from that of all models. We can verify this by comparing the numerical solution for $\Lambda$CDM to our analytical one. This comparison first requires to adjust the normalisation of the solution (4.24) as follows: we rewrite $X$ and $T$ as
@@
Figure 16: $a(t)$ in terms of $H_{{\rm today}}\,t$ (denoted $H_{0}\,t$ on the graphs) given by the analytical solution for the radiation - matter phase. Figure 16(a) gives this curve, analogous to that of Figure 15, but here in terms of time: we see that the red solution very quickly becomes the relevant one. In Figure 16(b), where we set $t_{{\rm today}}=0$, we compare this analytical solution (mostly red) to the numerical solution (blue) for $\Lambda$CDM. We verify the good match in the early universe (until dark matter domination), confirming in particular the normalisations discussed in the main text.
@@ 4.3.2 Analytic solution $\varphi^{i}(t)$
To solve $E^{i}=0$ and find the analytical solution for $\varphi^{i}(t)$, helpful hints are given by the numerical solutions. We depict in Figure 17 for our two quintessence examples the evolution of the kinetic energy and the potential, as well as that of the three different terms in $E^{i}$.
@@
Figure 17: Evolution of $V$, the kinetic energy $\rho_{{\rm kin}}=\frac{1}{2}\dot{\varphi}^{2}$, as well as of the different terms in the equation $E^{i}=0$, here for a single canonical field, namely $\partial_{\varphi}V$, $3H\dot{\varphi}$ and $\ddot{\varphi}$, in terms of $N$. This is depicted for our exponential and hilltop quintessence examples. For the latter, the quantities depicted are strictly speaking those divided by $H_{0}^{2}$, due to the formalism used for this example (see Section 2.2). The early times evolutions of $V$ and $\partial_{\varphi}V$ are related to the initial condition there being $\dot{\varphi}<0$ in our examples (see footnote 3).
@@
In this second part, the dynamics of the scalar fields is not only due to friction but also to the potential force. All three terms in $E^{i}=0$ are now relevant: we verify in Figure 17 that none can be neglected. This makes the equation more difficult to solve. Interestingly, we also note there that all three terms appear to be constant during most of this second part (and beyond). This is an important hint to find an analytical solution and we will come back to it.
@@
We have obtained an analytical solution for $\varphi^{i}(t)$ for the radiation - matter phase, that we separated in two parts. In the next subsection, we will specify better the normalisations of the correction terms in this solution, and make use of this result to compute the field variation during this phase. We will also use our physical understanding of this solution to evaluate the moment at which the potential becomes non-negligible, compared to the kinetic energy.
@@ 4.3.3 Applications: frozen fields ($\Delta\varphi$) and $w_{\varphi}$ transition
Having obtained the analytical solutions for $(a(t),\varphi^{i}(t))$ during the radiation - matter phase, we now want to make use of it to answer two questions: to what extent can we consider the scalar fields as “frozen” during this phase, and when is the transition moment at which $w_{\varphi}$ goes from $+1$ to $-1$?
@@
A more accurate estimate for $\Delta\varphi^{i}$ can be obtained from the analytical solution in (4.29) and (4.31). One unknown in these expressions are the normalisations in the time dependent pieces. A simplifying choice is to pick for $a(t)$ that of $P_{r}$ for the first part, and that of $P_{m}$ for the second part. This is justified by Figure 18, where the complete solution for $a(t)$ seems to be well approximated by these two successive scalings. From these choices, we obtain the following expressions
@@
This shows a remarkably good match with our previous analytical results. Note that above, we not even subtracted the evaluation at $N_{\varphi\,{\rm min}}$, and we mentioned it should be very small; this could improve the match of the last digits. This very good match confirms the parametrisation used for $a(t)$, the analytical solutions obtained, and the normalisation choices made, all resulting in the field expressions (4.33) and (4.34).
@@
The above analytical solution and our understanding of it allow us to determine $N_{{\rm kin}V}$. We explained that at first, the potential contributions were negligible, as during kination, and the kinetic energy was dominant but decreasing. Later, as can be seen in Figure 17(a) and 17(c), the kinetic energy becomes subdominant to the potential. This evolution is in line with the observed transition.
@@
We find a perfect match with the numerical values! This is a further confirmation of our analytical solutions.
@@ 4.4.1 Analytic solutions: attempts
As illustrated by this discussion, the dynamics in this last phase becomes at some point ($N\sim-1.5$) much more complicated to describe, as $\rho_{m}$, $V$ and then $\rho_{{\rm kin}}$ become all non-negligible. $F_{1}=0$ and $E^{i}=0$ become two coupled equations where no term can be neglected, and all are non-constant. We illustrate this in Figure 19. Finding an analytical solution is then much more difficult. In addition, the details of model, in particular the potential $V$, become very relevant, while our results were so far model-independent. Indeed, physically, the dynamics are such that the fields are rolling down the potential and gaining speed, while still facing (a not so high) Hubble friction: see Figure 17(a) and 17(c). The details of $V$ certainly matter in such dynamics.
@@
There is one case where we can still provide an analytical solution: $\Lambda$CDM. This result is well-known but let us reproduce it here for completeness (see [33] for a general solution in terms of a Jacobi elliptic function, in case of radiation, matter, curvature and cosmological constant). We take that $V=\Lambda=\rho_{\Lambda}>0$ and $\varphi^{i}$ are constant, so $E^{i}=0$ is automatically satisfied, and we are left to solve $F_{1}=0$. We expect to find a solution interpolating between $P_{m}$ and $P_{\Lambda}$. We can neglect radiation (due to its scaling in $a$), which amounts to set $\rho_{r0}=0$, and we restrict ourselves to the case without curvature, $k=0$. Then, $F_{1}=0$ gets rewritten as
@@
with $\Omega_{m}=0.315=1-\Omega_{\Lambda}$. This result is very close to the numerical value obtained in Figure 4. Note that the latter includes in addition the radiation - matter phase; it would be interesting to use the analytical results of Section 4.3.1 to evaluate the (very short) duration of that earlier phase.
Beyond this example, for quintessence models, it is difficult, as explained, to find an analytical solution. A useful parametrisation in the previous phase has been to consider $a(t)\sim t^{n_{a}}$, giving $H=n_{a}\,t^{-1}$ and $\ln H={\rm constant}-\frac{N}{n_{a}}$. Then, if the plot of $\ln H$ is a line, one can infer the value of $n_{a}$: see Figure 18. Focusing here on this last phase, we get the corresponding plots in Figure 20. For $\Lambda$CDM and hilltop quintessence, we do not get a line for $\ln H$, which discards this convenient parametrisation.
@@
In the case of exponential quintessence, one is close to getting a line around today’s universe, with $n_{a}\approx 1.04$. Further attempts using this observation are however inconclusive in getting an analytical solution. With the general study of exponential quintessence [15], we know that in our example, the asymptotic fixed point is $P_{\phi}$, giving a power law with $n_{a}=\frac{2}{3}$. This does not help in getting an analytical transient solution around today. We refrain from further attempts in getting analytical quintessence solutions in this last phase.
@@ 4.4.2 Integral results: $\Delta\varphi$ and $1+w_{\varphi}$
We did not obtain an analytical quintessence solution for the matter - dark energy phase, but we may still get interesting results thanks to integrals of relevant quantities over this period. In this subsubsection, we obtain one such results on the field distance $\Delta\varphi$ and one on the integral of $1+w_{\varphi}$.
@@
In case of canonically normalised fields, we obtain for each field the displacement $\Delta\varphi^{i}\leq\Delta\varphi$, with an equality for a single field. We leave aside the possible contribution of $V$ to the field distance, captured e.g. by the recently proposed generalised definitions of [34, 35]. In view of possible corrections to the effective theory that we use, it is important to verify whether the field distance remains sub-Planckian, meaning $\Delta\varphi\leq 1$. While its precise value is model and solution dependent, we aim here at getting an upper bound for it.
@@
The precise evolution of $\Omega_{\rm kin}$ is model dependent, but it has some universal features. First $\Omega_{\rm kin}$ starts close to zero, since we know it is negligible at and after $N_{m}$, until a schematic $N=N_{k}$, where typically $N_{k}\geq-1.5$. Second, $\Omega_{\rm kin}$ is growing at $N_{m}$ and after it: in our examples in Figure 19, it does so until today, where it reaches its value $\Omega_{\rm kin0}$. In the following we will assume such dynamics, namely a continuous growth from $N_{m}$ until today; note a difference with the evolution of $\rho_{\rm kin}$ in Figure 17. In case an example is found where $\Omega_{\rm kin}$ does not strictly grows during the matter - dark energy phase, we believe the difference with our treatment would be minor.
@@
We can refine this upper bound as follows. The above computation amounts to take a constant $\Omega_{\rm kin}$ over $N_{k}<N<0$, but we considered that it is growing from 00 to $\Omega_{\rm kin0}$ over this period. To capture the effect of this growth in the integral, let us consider as a very first approximation a linear growth. We take the simple parametrisation depicted in Figure 21, which amounts to
@@
for $N_{k}\geq-1.5$. This evaluation essentially gives a sub-Planckian field distance. Note that we would have obtained an even tighter bound if we had used in (4.50) the actual $w_{{\rm eff}0}$ value instead of $-1/3$.
@@
We recall that the right-hand side is the difference in maxima moments of $\Omega_{m}$ between $\Lambda$CDM and quintessence models, given identical $\Omega_{n0}$ today: we depict this difference in Figure 22(b). These $N_{m}$ values are given in Table 4 for our examples. Consistently, the difference vanishes for a cosmological constant. We also recall from the discussion around (3.27) that this difference is positive for quintessence due to Hubble friction; this is also expected from the left-hand side.
The analytical integral relation (4.60) between these two quantities is illustrated in Figure 22. It provides an estimate of the integral on the left-hand side.
@@
a result obtained for instance in [15, (3.13)] for a single field exponential quintessence, and shown here in multifield full generality. Through standard approximations, a similar integration, and making use of the results of Section 3.4, one reaches the same relation (4.60). In the following, we will investigate the implications of this relation, thanks to a specific parametrisation of $w_{\varphi}$.

### v2

@@ 4.3.1 Analytic solution $a(t)$
We turn to the radiation - matter phase that lasts between $N_{r}$ and $N_{m}$, as defined in (3.17). The phase starts with radiation domination and finishes with matter domination, in agreement with the scaling in $a$ of $\rho_{r}$ and $\rho_{m}$. We thus expect $P_{r}$ and $P_{m}$ to provide asymptotic expressions of a complete analytic solution for this whole phase, at small and large times respectively. We are interested in finding such a complete solution, interpolating between these two solutions.
@@
where $t_{*}$ captures the possible integration constant. This solution has been known for a long time: see [35, 34], and [36] with curvature, or [37, App. B] for a recent account.
@@
Figure 15: $X(T)$, standing up to normalisation for $a(t)$, as the analytic solution for the radiation - matter phase. The three colors of the curve correspond to the three analytic parts of the solution given in (4.24). The continuity of the solution is manifest.
@@
The analytic solution (4.24) should in very good approximation reproduce $a(t)$ during the radiation - matter phase, and in particular, reproduce the $\Lambda$CDM solution from $t=0$ until matter domination; in the later phase of dark energy domination, the analytic solution should differ from that of all models. We can verify this by comparing the numerical solution for $\Lambda$CDM to our analytic one. This comparison first requires to adjust the normalisation of the solution (4.24) as follows: we rewrite $X$ and $T$ as
@@
Figure 16: $a(t)$ in terms of $H_{{\rm today}}\,t$ (denoted $H_{0}\,t$ on the graphs) given by the analytic solution for the radiation - matter phase. Figure 16(a) gives this curve, analogous to that of Figure 15, but here in terms of time: we see that the red solution very quickly becomes the relevant one. In Figure 16(b), where we set $t_{{\rm today}}=0$, we compare this analytic solution (mostly red) to the numerical solution (blue) for $\Lambda$CDM. We verify the good match in the early universe (until dark matter domination), confirming in particular the normalisations discussed in the main text.
This properly normalised expression also allows to evaluate the duration of the radiation - matter phase. Strictly speaking, the phase starts at $N=N_{r}$ and ends at $N=N_{m}$, corresponding with the above radiation - matter equality normalisation, to $X_{r}=e^{N_{r}-N_{rm}}$ and $X_{m}=e^{N_{m}-N_{rm}}$. In good approximation, the former is $X_{r}\approx 0$, meaning that the phase can be considered here to start at $T=t=0$. Its duration is thus given by the time $t_{m}$ corresponding to $X_{m}$, namely
$H_{{\rm today}}\,t_{m}=\frac{2}{3}\,\Omega_{m0}^{-\frac{1}{2}}\,\frac{H_{{\rm today}}}{H_{0}}\,\left(2+(X_{m}-2)\sqrt{X_{m}+1}\right)\ ,$ (4.27)
with $\Omega_{m0}\approx 0.5000$, $H_{0}/H_{{\rm today}}=140236$ and $N_{rm}=-8.0548$. The result depends on $N_{m}$: with Table 4, we get the following durations for the radiation - matter phase
$\Lambda{\rm CDM}:\quad$ $H_{{\rm today}}\,t_{m}\approx 0.0285\ ,$
[…]
It is worth comparing these values to the numerical estimations of the age of the universe for each model in Section 3. In particular, analogously to the computation around (3.3), we get for $\Lambda$CDM that $t_{m}\approx 414\cdot 10^{6}\,{\rm yrs}$.
@@ 4.3.2 Analytic solution $\varphi^{i}(t)$
To solve $E^{i}=0$ and find the analytic solution for $\varphi^{i}(t)$, helpful hints are given by the numerical solutions. We depict in Figure 17 for our two quintessence examples the evolution of the kinetic energy and the potential, as well as that of the three different terms in $E^{i}$.
@@
Figure 17: Evolution of $V$, the kinetic energy $\rho_{{\rm kin}}=\frac{1}{2}\dot{\varphi}^{2}$, as well as of the different terms in the equation $E^{i}=0$, here for a single canonical field, namely $\partial_{\varphi}V$, $3H\dot{\varphi}$ and $\ddot{\varphi}$, in terms of $N$. This is depicted for our exponential and hilltop quintessence examples. Strictly speaking, the quantities depicted are those divided by $H_{0}^{2}$, due to the formalism used in these examples (see Section 2.2). The early times evolutions of $V$ and $\partial_{\varphi}V$ are related to the initial condition there being $\dot{\varphi}<0$ in our examples (see footnote 3).
@@
In this second part, the dynamics of the scalar fields is not only due to friction but also to the potential force. All three terms in $E^{i}=0$ are now relevant: we verify in Figure 17 that none can be neglected. This makes the equation more difficult to solve. Interestingly, we also note there that all three terms appear to be constant during most of this second part (and beyond). This is an important hint to find an analytic solution and we will come back to it.
@@
We have obtained an analytic solution for $\varphi^{i}(t)$ for the radiation - matter phase, that we separated in two parts. In the next subsection, we will specify better the normalisations of the correction terms in this solution, and make use of this result to compute the field variation during this phase. We will also use our physical understanding of this solution to evaluate the moment at which the potential becomes non-negligible, compared to the kinetic energy.
@@ 4.3.3 Applications: frozen fields ($\Delta\varphi$) and $w_{\varphi}$ transition
Having obtained the analytic solutions for $(a(t),\varphi^{i}(t))$ during the radiation - matter phase, we now want to make use of it to answer two questions: to what extent can we consider the scalar fields as “frozen” during this phase, and when is the transition moment at which $w_{\varphi}$ goes from $+1$ to $-1$?
@@
A more accurate estimate for $\Delta\varphi^{i}$ can be obtained from the analytic solution in (4.31) and (4.33). One unknown in these expressions are the normalisations in the time dependent pieces. A simplifying choice is to pick for $a(t)$ that of $P_{r}$ for the first part, and that of $P_{m}$ for the second part. This is justified by Figure 18, where the complete solution for $a(t)$ seems to be well approximated by these two successive scalings. From these choices, we obtain the following expressions
@@
This shows a remarkably good match with our previous analytical results. Note that above, we not even subtracted the evaluation at $N_{\varphi\,{\rm min}}$, and we mentioned it should be very small; this could improve the match of the last digits. This very good match confirms the parametrisation used for $a(t)$, the analytic solutions obtained, and the normalisation choices made, all resulting in the field expressions (4.35) and (4.36).
@@
The above analytic solution and our understanding of it allow us to determine $N_{{\rm kin}V}$. We explained that at first, the potential contributions were negligible, as during kination, and the kinetic energy was dominant but decreasing. Later, as can be seen in Figure 17(a) and 17(c), the kinetic energy becomes subdominant to the potential. This evolution is in line with the observed transition.
@@
We find a perfect match with the numerical values! This is a further confirmation of our analytic solutions.
@@ 4.4.1 Analytic solutions: attempts
As illustrated by this discussion, the dynamics in this last phase becomes at some point ($N\sim-1.5$) much more complicated to describe, as $\rho_{m}$, $V$ and then $\rho_{{\rm kin}}$ become all non-negligible. $F_{1}=0$ and $E^{i}=0$ become two coupled equations where no term can be neglected, and all are non-constant. We illustrate this in Figure 19. Finding an analytic solution is then much more difficult. In addition, the details of model, in particular the potential $V$, become very relevant, while our results were so far model-independent. Indeed, physically, the dynamics are such that the fields are rolling down the potential and gaining speed, while still facing (a not so high) Hubble friction: see Figure 17(a) and 17(c). The details of $V$ certainly matter in such dynamics.
@@
There is one case where we can still provide an analytic solution: $\Lambda$CDM. This result is well-known but let us reproduce it here for completeness (see [38] for a general solution in terms of a Jacobi elliptic function, in case of radiation, matter, curvature and cosmological constant). We take that $V=\Lambda=\rho_{\Lambda}>0$ and $\varphi^{i}$ are constant, so $E^{i}=0$ is automatically satisfied, and we are left to solve $F_{1}=0$. We expect to find a solution interpolating between $P_{m}$ and $P_{\Lambda}$. We can neglect radiation (due to its scaling in $a$), which amounts to set $\rho_{r0}=0$, and we restrict ourselves to the case without curvature, $k=0$. Then, $F_{1}=0$ gets rewritten as
@@
with $\Omega_{m}=0.315=1-\Omega_{\Lambda}$. This result is very close to the numerical value obtained in Figure 4. Cutting the above solution at the beginning of the matter - dark energy phase, namely at $N=N_{m}$, meaning removing the corresponding time from (4.49), and adding the duration the earlier radiation - matter phase computed in (4.28), we get even closer to the numerical value.
Beyond this example, for quintessence models, it is difficult, as explained, to find an analytic solution. A useful parametrisation in the previous phase has been to consider $a(t)\sim t^{n_{a}}$, giving $H=n_{a}\,t^{-1}$ and $\ln H={\rm constant}-\frac{N}{n_{a}}$. Then, if the plot of $\ln H$ is a line, one can infer the value of $n_{a}$: see Figure 18. Focusing here on this last phase, we get the corresponding plots in Figure 20. For $\Lambda$CDM and hilltop quintessence, we do not get a line for $\ln H$, which discards this convenient parametrisation.
@@
In the case of exponential quintessence, one is close to getting a line around today’s universe, with $n_{a}\approx 1.04$. Further attempts using this observation are however inconclusive in getting an analytic solution. With the general study of exponential quintessence [17], we know that in our example, the asymptotic fixed point is $P_{\phi}$, giving a power law with $n_{a}=\frac{2}{3}$. This does not help in getting an analytic transient solution around today. We refrain from further attempts in getting analytic quintessence solutions in this last phase.
@@ 4.4.2 Integral results: $\Delta\varphi$ and $1+w_{\varphi}$
We did not obtain an analytic quintessence solution for the matter - dark energy phase, but we may still get interesting results thanks to integrals of relevant quantities over this period. In this subsubsection, we obtain one such results on the field distance $\Delta\varphi$ and one on the integral of $1+w_{\varphi}$.
@@
In case of canonically normalised fields, we obtain for each field the displacement $\Delta\varphi^{i}\leq\Delta\varphi$, with an equality for a single field. We leave aside the possible contribution of $V$ to the field distance, captured e.g. by the recently proposed generalised definitions of [39, 40]. In view of possible corrections to the effective theory that we use, it is important to verify whether the field distance remains sub-Planckian (see e.g. the refined swampland distance conjecture [29]), meaning $\Delta\varphi\leq 1$. While its precise value is model and solution dependent, we aim here at getting an upper bound for it.
@@
The precise evolution of $\Omega_{\rm kin}$ is model dependent, but it has some universal features. First $\Omega_{\rm kin}$ starts close to zero, since we know it is negligible at and after $N_{m}$, until a schematic $N=N_{k}$, where typically $N_{k}\geq-1.5$. Second, $\Omega_{\rm kin}$ is growing at $N_{m}$ and after it: in our examples in Figure 19, it does so until today, where it reaches its value $\Omega_{\rm kin0}$. This continuous growth from $N_{m}$ until today can be understood from the expression $\Omega_{\rm kin}=\frac{1}{2}(1+w_{\varphi})\Omega_{\varphi}$. By definition of thawing models, and as will be discussed below, $w_{\varphi}$ is typically growing over this period of time. In addition, $\Omega_{\varphi}$ is also growing, as will be clear from (4.63) with $\Omega_{k}=0$ that gives $\partial_{N}\Omega_{\varphi}>0$. This results in a growth of $\Omega_{\rm kin}$ from $N_{m}$ until today, that will be assumed in the following; note a difference with the evolution of $\rho_{\rm kin}$, in e.g. Figure 17.
@@
We can refine this upper bound as follows. The above computation amounts to take a constant $\Omega_{\rm kin}$ over $N_{k}<N<0$, but we argued that it is growing from 00 to $\Omega_{\rm kin0}$ over this period. To capture the effect of this growth in the integral, let us consider as a very first approximation a linear growth. We take the simple parametrisation depicted in Figure 21, which amounts to
@@
for $N_{k}\geq-1.5$. This evaluation essentially gives a sub-Planckian field distance. Note that one can obtain a tighter bound depending on $w_{\varphi 0}$ by using the expression $\Omega_{\rm kin}=\frac{1}{2}(1+w_{\varphi})\Omega_{\varphi}$: this would give $\Delta\varphi\leq\frac{2}{3}\Delta N\,\sqrt{3\Omega_{\varphi 0}+3w_{\varphi 0}\Omega_{\varphi 0}}$, which is smaller than the above since $w_{\varphi 0}\Omega_{\varphi 0}<-1/3$ in an accelerating universe.
@@
We recall that the right-hand side is the difference in maxima moments of $\Omega_{m}$ between $\Lambda$CDM and quintessence models, given identical $\Omega_{n0}$ today: we depict this difference in Figure 22(b) (a similar figure can be found in [41], with related discussions). These $N_{m}$ values are given in Table 4 for our examples. Consistently, the difference vanishes for a cosmological constant. We also recall from the discussion around (3.27) that this difference is positive for quintessence due to Hubble friction; this is also expected from the left-hand side.
The analytic integral relation (4.62) between these two quantities is illustrated in Figure 22. Around (5.1), we discuss an analogous relation between a similar integral and the moments of matter - dark energy equality, the difference of which can also be seen in Figure 22(b).
@@
a result obtained for instance in [17, (3.13)] for a single field exponential quintessence, and shown here in multifield full generality. Through standard approximations, a similar integration, and making use of the results of Section 3.4, one reaches the same relation (4.62). In the following, we will investigate the implications of this relation, thanks to a specific parametrisation of $w_{\varphi}$.